# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 7 One side flagness in $H_{3}$
Let $\Lambda$ be the Dynkin diagram of type $H_{3}$ with vertex set $\{a,b,c\}$ such that $m_{ab}=5$, $m_{bc}=3$. Let $\Delta$ be the Artin complex of the Artin group $A_{\Lambda}$ with the order on its vertex induced from $a<b<c$ (see Definition 3.12). Then $\Delta$ is downward flag.
@@
We denote the sub-segments of this path by $\{P_{i}\}_{i=1}^{6}$. Let $\mathsf{C}$ be the Coxeter complex of type $H_{3}$. Let $\pi:\Delta\to\mathsf{C}$ be the map induced by the action of the pure Artin group on $\Delta$.
Case 1: the $\pi$-image of the 6-cycle $\omega$ is a single edge $\bar{x}\bar{y}$ in $\mathsf{C}$. Let $C_{\bar{x}}$ (resp. $C_{\bar{y}}$) be the 2-cell in $\Sigma$ dual to $\bar{x}$ (resp. $\bar{y}$). Let $\widehat{C}_{\bar{x}}$ and $\widehat{C}_{\bar{y}}$ be the associated standard subcomplexes of $\widehat{\Sigma}$. Up to replacing $w=\Pi_{i=1}^{6}w_{i}$ by an equivalent word (in the sense of Definition 3.6), we can assume $P_{i}$ is a loop in $\widehat{C}_{\bar{x}}$ (resp. $\widehat{C}_{\bar{y}}$) for $i$ odd (resp. for $i$ even). Let $\widehat{Z}$ be the complex defined before Lemma 6.5. Recall from Section 6.1 that there is an embedding $\widehat{Z}\to\widehat{\Sigma}$. Thus we can also view $\widehat{Z}$ as a subcomplex of $\widehat{\Sigma}$. Recall that edges of $\widehat{\Sigma}$ are oriented and labeled by $\{a,b,c\}$ (Definition 2.11), this also gives label and orientation of edges in $\widehat{Z}$ via the inclusion of $\widehat{Z}$ into $\widehat{\Sigma}$. Up to a symmetry of $\widehat{\Sigma}$, we can assume
@@
Let $\mathbb{D}\to\mathbb{U}$ be a minimal area singular disk diagram (cf. Section 2.8) for the loop $\omega_{\mathbb{U}}$. We will slightly abuse notation and use $z_{i}$ to denote the point in the boundary cycle of $\mathbb{D}$ mapping to $z_{i}\in\mathbb{U}$. The metric on $\mathbb{U}$ induces a metric on $\mathbb{D}$, which is CAT$(0)$. For three points $x,y,z$ in a CAT$(0)$ space, we use $\angle_{y}(x,z)$ to denote the Alexandrov between the geodesic segments $\overline{yx}$ and $\overline{yz}$ (cf. [BH99, Chapter II.3.1]).
@@
For each vertex $z_{i}$ on the boundary cycle $\omega_{\mathbb{U}}$ of $\mathbb{D}$, the interval angle at $z_{i}$ is the quantity $\alpha(z_{i})$ defined in Section 2.8. By the combinatorial Gauss-Bonnet formula in Section 2.8, least four internal angles of $\omega_{\mathbb{U}}$ in $\mathbb{D}$ is $\pi/2$.
@@
    If $\angle_{z_{2}}(z_{1},z_{31})\geq 3\pi/4$ and $\angle_{z_{4}}(z_{5},z_{33})\geq 3\pi/4$ in $\mathbb{D}$, then from the Gauss-Bonnet formula (2.16) we know all the other angles at the vertices of $\omega_{\mathbb{U}}$ in $\mathbb{D}$ has to be $\pi/2$. As $\mathbb{D}$ is made of right-angled isosceles triangles, by consider the sequence of right-angled turns in the edge path $z_{4}\to z_{5}\to z_{6}\to z_{1}\to z_{2}$, we must have $z_{2}=z_{4}$, which is impossible as $z_{2}\to z_{31}\to z_{32}\to z_{33}\to z_{4}$ is a geodesic in $\mathbb{U}$.
@@
    If $\angle_{z_{2}}(z_{1},z_{31})=\pi/4$ and $\angle_{z_{4}}(z_{5},z_{33})\geq 3\pi/4$ in $\mathbb{D}$, then $z_{2},z_{31},z_{1}$ form the vertices of a triangle with right-angled at $z_{31}$. Thus $z_{1}$ is adjacent to $z_{32}$. Now we are reduced to consider a new cycle $\omega^{\prime}_{\mathbb{U}}$ of form $z_{1}\to z_{32}\to z_{33}\to z_{4}\to z_{5}\to z_{6}\to z_{1}$ and a disk diagram $\mathbb{D}^{\prime}\subset\mathbb{D}$ bounded by $\omega^{\prime}_{\mathbb{U}}$. Note that $\angle_{z_{32}}(z_{1},z_{33})=3\pi/4$ in $\mathbb{D}^{\prime}$. Thus by Gauss-Bonnet formula, all other angles at the vertices of $\omega^{\prime}_{\mathbb{U}}$ in $\mathbb{D}^{\prime}$ is $\pi/2$. We deduce as before that $z_{32}=z_{4}$, which is impossible as $z_{32}\to z_{33}\to z_{4}$ is a geodesic in $\mathbb{U}$.
@@
$\widehat{\mathsf{X}},\widehat{\mathsf{X}}^{\prime}$, $\{\widehat{\mathsf{X}}_{i}\}_{i=1}^{4}$, $\widehat{\mathsf{X}}_{ij}$, $\widehat{\mathsf{Y}}$ and $c:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{K}}$
be as in Section 6.2. We view $\widehat{\mathsf{X}}$ as a subcomplex of $\widehat{\Sigma}_{\mathcal{K}}$. Let $\hat{c}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{K}}$ be the map induced by $c$. We will assume
$\hat{c}(\widehat{C}_{1})=\widehat{\mathsf{X}}_{22}$, $\hat{c}(\widehat{C}_{2})=\widehat{\mathsf{X}}_{31}$, $\hat{c}(\widehat{C}_{3})=\widehat{\mathsf{X}}_{42}$ and $\hat{c}(\widehat{C}_{4})=\widehat{\mathsf{X}}_{33}$.
Then $\hat{c}(P)\subset\widehat{\mathsf{X}}^{\prime}$ and $\hat{c}(P)$ is null-homotopic in $\widehat{\mathsf{X}}^{\prime}$ by Lemma 2.13 and Lemma 6.8. Let $W_{234}$ be the Coxeter group of type $B_{3}$ with its generating set $S=\{r,s,t\}$ such that $m_{rs}=4$ and $m_{st}=3$. Let $\widehat{\Sigma}_{234}$ and $\hat{e}:\widehat{\mathsf{X}}^{\prime}\to\widehat{\Sigma}_{234}$ be as in Section 6.2. Let $\Delta_{234}$ be the Artin complex of the Artin group of type $B_{3}$. Then $Q=\hat{e}\circ\hat{c}(P)$ is null-homotopic in $\widehat{\Sigma}_{234}$. Let $\widetilde{\Sigma}_{234}$ be the universal cover of $\widehat{\Sigma}_{234}$. We define a subcomplex of type $\hat{r}$ in $\widetilde{\Sigma}_{234}$ to be a lift of a standard subcomplex of type $\hat{r}$ in $\widehat{\Sigma}_{234}$.
Let $Q_{i}=\hat{e}\circ\hat{c}(P_{i})$. Let $\widetilde{Q}$ be a lift of $Q$ to $\widetilde{\Sigma}_{234}$. Let $\widehat{C}^{\prime}_{i}=\hat{e}\circ\hat{c}(\widehat{C}_{i})$, and $\widehat{C}^{\prime}_{ij}=\hat{e}\circ\hat{c}(\widehat{C}_{ij})$. For $i=1,3,5$, let $\widetilde{\Phi}_{i}$ be the subcomplex of type $\hat{t}$ that contains $\widetilde{Q}_{i}$, and let $y^{\prime}_{i}$ be the vertex in $\Delta_{234}$ corresponding to $\widetilde{\Phi}_{i}$. For $i=2,4,6$, let $\widetilde{\Phi}_{i}$ be the subcomplex of type $\hat{r}$ that contains $\widetilde{Q}_{i}$, and let $x^{\prime}_{i}$ be the vertex in $\Delta_{234}$ corresponding to $\widetilde{\Phi}_{i}$. Let $\omega^{\prime}$ be the following 6-cycle in $\Delta_{234}$:
@@
We now show $\omega^{\prime}$ is locally embedded. Clearly $y^{\prime}_{1}\neq y^{\prime}_{i}$ for $i=2,3$ as $\widetilde{\Phi}_{1}$ and $\widetilde{\Phi}_{i}$ are mapped to different standard subcomplexes in $\widehat{\Sigma}_{234}$. Similar reason yields $x^{\prime}_{2}\neq x^{\prime}_{i}$ for $i=1,3$. From the description of $P_{1}$ and $P_{4}$, we know $\hat{c}$ does not collapse any edges in $P_{1}$ and $P_{4}$. Thus the loop $Q_{1}$ represents an element in $\pi_{1}(\widehat{C}^{\prime}_{1})$ which is not contained in $\pi_{1}(\widehat{C}^{\prime}_{1}\cap\widehat{C}^{\prime}_{2})$, which implies $x^{\prime}_{1}\neq x^{\prime}_{3}$. Similarly $y^{\prime}_{2}\neq y^{\prime}_{3}$.
@@
Then $\widetilde{Q}^{\prime}_{1}$ is homotopic to $\widetilde{Q}_{1}$ rel endpoints in $\widetilde{\Phi}_{1}$. Let $\Phi_{i}$ be the image of $\widetilde{\Phi}_{i}$ under $\widetilde{\Sigma}_{234}\to\widehat{\Sigma}_{234}$. We define $\Phi_{w}$ and $Q^{\prime}_{3}$ similarly. Let $\bar{w},\bar{x}^{\prime}_{i}$ be the image of $w,x^{\prime}_{i}$ under $\Delta_{234}\to\mathsf{C}_{234}$ respectively. Then $\bar{w}$ is adjacent $\bar{y}^{\prime}_{1},\bar{y}^{\prime}_{2}$ and $\bar{x}^{\prime}_{1}$. This determines the position of $\bar{w}$. In particular,
@@
Thus $Q^{\prime}_{1}\subset\widehat{C}^{\prime}_{13}\cup\widehat{C}^{\prime}_{12}$. Let $Q_{1i}=\hat{e}\circ\hat{c}(P_{1i})$. Then $Q_{11}\cup Q_{13}\subset\widehat{C}^{\prime}_{12}$ and $Q_{12}\subset\widehat{C}^{\prime}_{14}$. Consider the retraction $r:\widehat{C}^{\prime}_{1}\to\widehat{C}^{\prime}_{14}$. As $Q_{1}$ and $Q^{\prime}_{1}$ are homotopic rel endpoints in $\widehat{C}^{\prime}_{1}=\Phi_{1}$, we know $r(Q_{1})$ and $r(Q^{\prime}_{1})$ are homotopic rel endpoint in $\widehat{C}^{\prime}_{14}$. However, this is a contradiction as $r(Q^{\prime}_{1})$ is a point, and $r(Q_{1})=Q_{12}$ is null-homotopic as we are assuming $k_{2}\neq 0$ and $k_{5}\neq 0$. Thus $z\neq x^{\prime}_{1}$. The proof of $z\neq x^{\prime}_{3}$ is similar. ∎
@@
We only prove the claim for $w_{bc}$, as the other case is similar. Consider the 4-cycle $y^{\prime}_{1}x^{\prime}_{1}zx^{\prime}_{3}$, which is embedded by Claim 7.5. By Theorem 3.17, there exists a vertex $w^{\prime}\in\Delta_{234}$ of type $\hat{s}$ such that it is adjacent of the each of the vertices of this 4-cycle. Let $\bar{w}^{\prime}$ be the image of $w^{\prime}$ under $\Delta_{234}\to\mathsf{C}_{234}$. Then $\bar{w}^{\prime}$ is the unique vertex which is adjacent to $\bar{y}^{\prime}_{1},\bar{y}^{\prime}_{2}$ and $\bar{x}^{\prime}_{1}$. By the same argument as in the proof of Claim 2, we know $Q_{2}$ is homotopic rel endpoints in $\widehat{C}^{\prime}_{2}$ to a path $Q^{\prime}_{2}$ which is a concatenation of $Q^{\prime}_{21}\subset\widehat{C}^{\prime}_{21}$, $Q^{\prime}_{22}\subset\widehat{C}^{\prime}_{24}$ and $Q^{\prime}_{23}\subset\widehat{C}^{\prime}_{23}$. As $\hat{e}\circ\hat{c}$ map $\widehat{C}_{2}$ homeomorphically onto $\widehat{C}^{\prime}_{2}$, we know $P_{2}$ is homotopic rel endpoints in $\widehat{C}_{2}$ to the concatenation of a path in $\widehat{C}_{21}$, a path in $\widehat{C}_{24}$ and a path in $\widehat{C}_{23}$. Hence $w_{bc}=b^{*}c^{*}b^{*}$. ∎
@@
By Theorem 3.22, there is a vertex $z\in\Delta_{234}$ of type $\hat{t}$ such that it is adjacent to $y^{\prime}_{i}$ for $1\leq i\leq 3$. Let $\mathsf{C}_{234}$ be the Coxeter complex for $W_{234}$. Recall that there is a map $\Delta_{234}\to\mathsf{C}_{234}$ induced by the quotient of the action of the pure Artin group $PA_{234}$. Let $\bar{z}$ be the image of $z$ under the map $\Delta_{234}\to\mathbb{\mathsf{C}}_{234}$. We define $\bar{x}^{\prime}_{i}$ and $\bar{y}^{\prime}_{i}$ similarly. As $\bar{z}$ is adjacent to $\bar{y}^{\prime}_{1}$, $\bar{y}^{\prime}_{2}$ and $\bar{y}^{\prime}_{3}$, we know $\bar{x}^{\prime}_{1}=\bar{x}^{\prime}_{2}=\bar{x}^{\prime}_{3}=\bar{z}$.
@@
As $\bar{w}_{j}$ is adjacent to $\bar{z},\bar{y}^{\prime}_{j}$ and $\bar{y}^{\prime}_{j+1}$, this uniquely determine the position of $\bar{w}_{j}$ in $\mathsf{C}_{234}$. This implies that
@@
Case 5.2: $\bar{x}$ is of type $\hat{c}$. We write $\bar{y}_{i}\sim\bar{y}_{j}$ if they are adjacent to the same vertex of type $\hat{b}$ in $\mathsf{C}$. Up to a permutation of $\{\bar{y}_{i}\}_{i=1}^{3}$, there are only two possibilities, namely either $\bar{y}_{1}\sim\bar{y}_{2}$ and $\bar{y}_{2}\sim\bar{y}_{3}$, or $\bar{y}_{1}\sim\bar{y}_{3},\bar{y}_{1}\nsim\bar{y}_{2}$ and $y_{2}\nsim y_{3}$. We assume $P_{i}\subset\widehat{C}_{\bar{x}}$ for $i=1,3,5$, and $P_{2i}\subset\widehat{C}_{y_{i}}$ for $i=1,2,3$.
@@
Case 6: the $\pi$-image of $\omega$ is four edges. In this case, the only possibility is that $\pi(\omega)$ is a 4-cycle. Let $\bar{x}_{1},\bar{x}_{2},\bar{x}_{3},\bar{x}_{4}$ be consecutive vertices on this 4-cycle. Let $\widehat{C}_{i}=\widehat{C}_{\bar{x}_{i}}$ and $\widehat{C}_{ij}=\Pi_{\widehat{C}_{i}}(\widehat{C}_{j})$. Up to a cyclic permutation of the index and symmetries of $\mathsf{C}$, we are in one of the following situations:

### v3

@@ 7 One side flagness in $H_{3}$
Let $\Lambda$ be the Coxeter diagram of type $H_{3}$ with the set of nodes $\{a,b,c\}$ such that $m_{ab}=5$, $m_{bc}=3$. Let $\Delta$ be the Artin complex of the Artin group $A_{\Lambda}$ with the order on its vertex induced from $a<b<c$ (see Definition 2.15). Then $\Delta$ is downward flag.
@@
We need a procedure of converting a $n$-cycle in the Artin complex $\Delta=\Delta_{\Lambda}$ of $A_{\Lambda}$ to a concatenation of $n$ words in $A_{\Lambda}$ (cf. [Hua24a, Definition 6.14]). These $n$ words are well-defined up to an appropriate notion of equivalence.
Definition 7.2.
@@
In the following proof we will need a simple version of combinatorial Gauss-Bonnet formula. We refer to [Hua24a, Section 3.6] for a quick summary.
@@
We denote the sub-segments of this path by $\{P_{i}\}_{i=1}^{6}$. Let $\mathfrak{C}$ be the Coxeter complex of type $H_{3}$. Let $\pi:\Delta\to\mathfrak{C}$ be the map induced by the action of the pure Artin group on $\Delta$.
Case 1: the $\pi$-image of the 6-cycle $\omega$ is a single edge $\bar{x}\bar{y}$ in $\mathfrak{C}$. Let $C_{\bar{x}}$ (resp. $C_{\bar{y}}$) be the 2-cell in $\Sigma$ dual to $\bar{x}$ (resp. $\bar{y}$). Let $\widehat{C}_{\bar{x}}$ and $\widehat{C}_{\bar{y}}$ be the associated standard subcomplexes of $\widehat{\Sigma}$. Up to replacing $w=\Pi_{i=1}^{6}w_{i}$ by an equivalent word (in the sense of Definition 7.2), we can assume $P_{i}$ is a loop in $\widehat{C}_{\bar{x}}$ (resp. $\widehat{C}_{\bar{y}}$) for $i$ odd (resp. for $i$ even). Let $\widehat{Z}$ be the complex defined before Lemma 6.5. Recall from Section 6.1 that there is an embedding $\widehat{Z}\to\widehat{\Sigma}$. Thus we can also view $\widehat{Z}$ as a subcomplex of $\widehat{\Sigma}$. Recall that edges of $\widehat{\Sigma}$ are oriented and labeled by $\{a,b,c\}$ (Definition 4.8), this also gives label and orientation of edges in $\widehat{Z}$ via the inclusion of $\widehat{Z}$ into $\widehat{\Sigma}$. Up to a symmetry of $\widehat{\Sigma}$, we can assume
@@
Let $\mathbb{D}\to\mathbb{U}$ be a minimal area singular disk diagram ([Hua24a, Section 3.6]) for the loop $\omega_{\mathbb{U}}$. We will slightly abuse notation and use $z_{i}$ to denote the point in the boundary cycle of $\mathbb{D}$ mapping to $z_{i}\in\mathbb{U}$. The metric on $\mathbb{U}$ induces a metric on $\mathbb{D}$, which is CAT$(0)$. For three points $x,y,z$ in a CAT$(0)$ space, we use $\angle_{y}(x,z)$ to denote the Alexandrov between the geodesic segments $\overline{yx}$ and $\overline{yz}$ (cf. [BH99, Chapter II.3.1]).
@@
For each vertex $z_{i}$ on the boundary cycle $\omega_{\mathbb{U}}$ of $\mathbb{D}$, the interval angle at $z_{i}$ is the quantity $\alpha(z_{i})$ defined in [Hua24a, Section 3.6]. By the combinatorial Gauss-Bonnet formula in [Hua24a, Equation 3.17], at least four internal angles of $\omega_{\mathbb{U}}$ in $\mathbb{D}$ is $\pi/2$.
@@
    If $\angle_{z_{2}}(z_{1},z_{31})\geq 3\pi/4$ and $\angle_{z_{4}}(z_{5},z_{33})\geq 3\pi/4$ in $\mathbb{D}$, then from [Hua24a, Equation 3.17] we know all the other angles at the vertices of $\omega_{\mathbb{U}}$ in $\mathbb{D}$ has to be $\pi/2$. As $\mathbb{D}$ is made of right-angled isosceles triangles, by consider the sequence of right-angled turns in the edge path $z_{4}\to z_{5}\to z_{6}\to z_{1}\to z_{2}$, we must have $z_{2}=z_{4}$, which is impossible as $z_{2}\to z_{31}\to z_{32}\to z_{33}\to z_{4}$ is a geodesic in $\mathbb{U}$.
@@
    If $\angle_{z_{2}}(z_{1},z_{31})=\pi/4$ and $\angle_{z_{4}}(z_{5},z_{33})\geq 3\pi/4$ in $\mathbb{D}$, then $z_{2},z_{31},z_{1}$ form the vertices of a triangle with right-angled at $z_{31}$. Thus $z_{1}$ is adjacent to $z_{32}$. Now we are reduced to consider a new cycle $\omega^{\prime}_{\mathbb{U}}$ of form $z_{1}\to z_{32}\to z_{33}\to z_{4}\to z_{5}\to z_{6}\to z_{1}$ and a disk diagram $\mathbb{D}^{\prime}\subset\mathbb{D}$ bounded by $\omega^{\prime}_{\mathbb{U}}$. Note that $\angle_{z_{32}}(z_{1},z_{33})=3\pi/4$ in $\mathbb{D}^{\prime}$. Thus by [Hua24a, Equation 3.17], all other angles at the vertices of $\omega^{\prime}_{\mathbb{U}}$ in $\mathbb{D}^{\prime}$ is $\pi/2$. We deduce as before that $z_{32}=z_{4}$, which is impossible as $z_{32}\to z_{33}\to z_{4}$ is a geodesic in $\mathbb{U}$.
@@
$\widehat{\mathsf{X}},\widehat{\mathsf{X}}^{\prime}$, $\{\widehat{\mathsf{X}}_{i}\}_{i=1}^{4}$, $\widehat{\mathsf{X}}_{ij}$, $\widehat{\mathsf{Y}}$ and $\kappa:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{K}}$
be as in Section 6.2. We view $\widehat{\mathsf{X}}$ as a subcomplex of $\widehat{\Sigma}_{\mathcal{K}}$. Let $\hat{\kappa}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{K}}$ be the map induced by $\kappa$. We will assume
$\hat{\kappa}(\widehat{C}_{1})=\widehat{\mathsf{X}}_{22}$, $\hat{\kappa}(\widehat{C}_{2})=\widehat{\mathsf{X}}_{31}$, $\hat{\kappa}(\widehat{C}_{3})=\widehat{\mathsf{X}}_{42}$ and $\hat{\kappa}(\widehat{C}_{4})=\widehat{\mathsf{X}}_{33}$.
Then $\hat{\kappa}(P)\subset\widehat{\mathsf{X}}^{\prime}$ and $\hat{\kappa}(P)$ is null-homotopic in $\widehat{\mathsf{X}}^{\prime}$ by Lemma 4.11 and Lemma 6.8. Let $W_{234}$ be the Coxeter group of type $B_{3}$ with its generating set $S=\{r,s,t\}$ such that $m_{rs}=4$ and $m_{st}=3$. Let $\widehat{\Sigma}_{234}$ and $\hat{e}:\widehat{\mathsf{X}}^{\prime}\to\widehat{\Sigma}_{234}$ be as in Section 6.2. Let $\Delta_{234}$ be the Artin complex of the Artin group of type $B_{3}$. Then $Q=\hat{e}\circ\hat{\kappa}(P)$ is null-homotopic in $\widehat{\Sigma}_{234}$. Let $\widetilde{\Sigma}_{234}$ be the universal cover of $\widehat{\Sigma}_{234}$. We define a subcomplex of type $\hat{r}$ in $\widetilde{\Sigma}_{234}$ to be a lift of a standard subcomplex of type $\hat{r}$ in $\widehat{\Sigma}_{234}$.
Let $Q_{i}=\hat{e}\circ\hat{\kappa}(P_{i})$. Let $\widetilde{Q}$ be a lift of $Q$ to $\widetilde{\Sigma}_{234}$. Let $\widehat{C}^{\prime}_{i}=\hat{e}\circ\hat{\kappa}(\widehat{C}_{i})$, and $\widehat{C}^{\prime}_{ij}=\hat{e}\circ\hat{\kappa}(\widehat{C}_{ij})$. For $i=1,3,5$, let $\widetilde{\Phi}_{i}$ be the subcomplex of type $\hat{t}$ that contains $\widetilde{Q}_{i}$, and let $y^{\prime}_{i}$ be the vertex in $\Delta_{234}$ corresponding to $\widetilde{\Phi}_{i}$. For $i=2,4,6$, let $\widetilde{\Phi}_{i}$ be the subcomplex of type $\hat{r}$ that contains $\widetilde{Q}_{i}$, and let $x^{\prime}_{i}$ be the vertex in $\Delta_{234}$ corresponding to $\widetilde{\Phi}_{i}$. Let $\omega^{\prime}$ be the following 6-cycle in $\Delta_{234}$:
@@
We now show $\omega^{\prime}$ is locally embedded. Clearly $y^{\prime}_{1}\neq y^{\prime}_{i}$ for $i=2,3$ as $\widetilde{\Phi}_{1}$ and $\widetilde{\Phi}_{i}$ are mapped to different standard subcomplexes in $\widehat{\Sigma}_{234}$. Similar reason yields $x^{\prime}_{2}\neq x^{\prime}_{i}$ for $i=1,3$. From the description of $P_{1}$ and $P_{4}$, we know $\hat{\kappa}$ does not collapse any edges in $P_{1}$ and $P_{4}$. Thus the loop $Q_{1}$ represents an element in $\pi_{1}(\widehat{C}^{\prime}_{1})$ which is not contained in $\pi_{1}(\widehat{C}^{\prime}_{1}\cap\widehat{C}^{\prime}_{2})$, which implies $x^{\prime}_{1}\neq x^{\prime}_{3}$. Similarly $y^{\prime}_{2}\neq y^{\prime}_{3}$.
@@
Then $\widetilde{Q}^{\prime}_{1}$ is homotopic to $\widetilde{Q}_{1}$ rel endpoints in $\widetilde{\Phi}_{1}$. Let $\Phi_{i}$ be the image of $\widetilde{\Phi}_{i}$ under $\widetilde{\Sigma}_{234}\to\widehat{\Sigma}_{234}$. We define $\Phi_{w}$ and $Q^{\prime}_{3}$ similarly. Let $\bar{w},\bar{x}^{\prime}_{i}$ be the image of $w,x^{\prime}_{i}$ under $\Delta_{234}\to\mathfrak{C}_{234}$ respectively. Then $\bar{w}$ is adjacent $\bar{y}^{\prime}_{1},\bar{y}^{\prime}_{2}$ and $\bar{x}^{\prime}_{1}$. This determines the position of $\bar{w}$. In particular,
@@
Thus $Q^{\prime}_{1}\subset\widehat{C}^{\prime}_{13}\cup\widehat{C}^{\prime}_{12}$. Let $Q_{1i}=\hat{e}\circ\hat{\kappa}(P_{1i})$. Then $Q_{11}\cup Q_{13}\subset\widehat{C}^{\prime}_{12}$ and $Q_{12}\subset\widehat{C}^{\prime}_{14}$. Consider the retraction $r:\widehat{C}^{\prime}_{1}\to\widehat{C}^{\prime}_{14}$. As $Q_{1}$ and $Q^{\prime}_{1}$ are homotopic rel endpoints in $\widehat{C}^{\prime}_{1}=\Phi_{1}$, we know $r(Q_{1})$ and $r(Q^{\prime}_{1})$ are homotopic rel endpoint in $\widehat{C}^{\prime}_{14}$. However, this is a contradiction as $r(Q^{\prime}_{1})$ is a point, and $r(Q_{1})=Q_{12}$ is null-homotopic as we are assuming $k_{2}\neq 0$ and $k_{5}\neq 0$. Thus $z\neq x^{\prime}_{1}$. The proof of $z\neq x^{\prime}_{3}$ is similar. ∎
@@
We only prove the claim for $w_{bc}$, as the other case is similar. Consider the 4-cycle $y^{\prime}_{1}x^{\prime}_{1}zx^{\prime}_{3}$, which is embedded by Claim 7.6. By Theorem 2.35, there exists a vertex $w^{\prime}\in\Delta_{234}$ of type $\hat{s}$ such that it is adjacent of the each of the vertices of this 4-cycle. Let $\bar{w}^{\prime}$ be the image of $w^{\prime}$ under $\Delta_{234}\to\mathfrak{C}_{234}$. Then $\bar{w}^{\prime}$ is the unique vertex which is adjacent to $\bar{y}^{\prime}_{1},\bar{y}^{\prime}_{2}$ and $\bar{x}^{\prime}_{1}$. By the same argument as in the proof of Claim 2, we know $Q_{2}$ is homotopic rel endpoints in $\widehat{C}^{\prime}_{2}$ to a path $Q^{\prime}_{2}$ which is a concatenation of $Q^{\prime}_{21}\subset\widehat{C}^{\prime}_{21}$, $Q^{\prime}_{22}\subset\widehat{C}^{\prime}_{24}$ and $Q^{\prime}_{23}\subset\widehat{C}^{\prime}_{23}$. As $\hat{e}\circ\hat{\kappa}$ map $\widehat{C}_{2}$ homeomorphically onto $\widehat{C}^{\prime}_{2}$, we know $P_{2}$ is homotopic rel endpoints in $\widehat{C}_{2}$ to the concatenation of a path in $\widehat{C}_{21}$, a path in $\widehat{C}_{24}$ and a path in $\widehat{C}_{23}$. Hence $w_{bc}=b^{*}c^{*}b^{*}$. ∎
@@
By Theorem 2.37, there is a vertex $z\in\Delta_{234}$ of type $\hat{t}$ such that it is adjacent to $y^{\prime}_{i}$ for $1\leq i\leq 3$. Let $\mathfrak{C}_{234}$ be the Coxeter complex for $W_{234}$. Recall that there is a map $\Delta_{234}\to\mathfrak{C}_{234}$ induced by the quotient of the action of the pure Artin group $PA_{234}$. Let $\bar{z}$ be the image of $z$ under the map $\Delta_{234}\to\mathbb{\mathfrak{C}}_{234}$. We define $\bar{x}^{\prime}_{i}$ and $\bar{y}^{\prime}_{i}$ similarly. As $\bar{z}$ is adjacent to $\bar{y}^{\prime}_{1}$, $\bar{y}^{\prime}_{2}$ and $\bar{y}^{\prime}_{3}$, we know $\bar{x}^{\prime}_{1}=\bar{x}^{\prime}_{2}=\bar{x}^{\prime}_{3}=\bar{z}$.
@@
As $\bar{w}_{j}$ is adjacent to $\bar{z},\bar{y}^{\prime}_{j}$ and $\bar{y}^{\prime}_{j+1}$, this uniquely determine the position of $\bar{w}_{j}$ in $\mathfrak{C}_{234}$. This implies that
@@
Case 5.2: $\bar{x}$ is of type $\hat{c}$. We write $\bar{y}_{i}\sim\bar{y}_{j}$ if they are adjacent to the same vertex of type $\hat{b}$ in $\mathfrak{C}$. Up to a permutation of $\{\bar{y}_{i}\}_{i=1}^{3}$, there are only two possibilities, namely either $\bar{y}_{1}\sim\bar{y}_{2}$ and $\bar{y}_{2}\sim\bar{y}_{3}$, or $\bar{y}_{1}\sim\bar{y}_{3},\bar{y}_{1}\nsim\bar{y}_{2}$ and $y_{2}\nsim y_{3}$. We assume $P_{i}\subset\widehat{C}_{\bar{x}}$ for $i=1,3,5$, and $P_{2i}\subset\widehat{C}_{y_{i}}$ for $i=1,2,3$.
@@
Case 6: the $\pi$-image of $\omega$ is four edges. In this case, the only possibility is that $\pi(\omega)$ is a 4-cycle. Let $\bar{x}_{1},\bar{x}_{2},\bar{x}_{3},\bar{x}_{4}$ be consecutive vertices on this 4-cycle. Let $\widehat{C}_{i}=\widehat{C}_{\bar{x}_{i}}$ and $\widehat{C}_{ij}=\Pi_{\widehat{C}_{i}}(\widehat{C}_{j})$. Up to a cyclic permutation of the index and symmetries of $\mathfrak{C}$, we can assume $\bar{x}_{1}$ is of type $\hat{\hat{c}}$, and one of the following holds true: